### v1

# DiffHairCard: Auto Hair Card Extraction with Differentiable Rendering

Abstract.
Hair cards remain a widely used representation for hair modeling in real-time applications, offering a practical trade-off between visual fidelity, memory usage, and performance. However, generating high-quality hair card models remains a challenging and labor-intensive task. This work presents an automated pipeline for converting strand-based hair models into hair card models with a limited number of cards and textures while preserving the hairstyle appearance. Our key idea is a novel differentiable representation where each strand is encoded as a projected 2D curve in the texture space, which enables end-to-end optimization with differentiable rendering while respecting the structures of the hair geometry. Based on this representation, we develop a novel algorithm pipeline, where we first cluster hair strands into initial hair cards and project the strands into the texture space. We then conduct a two-stage optimization, where our first stage optimizes the orientation of each hair card separately, and after strand projection, our second stage conducts joint optimization over the entire hair card model for fine-tuning. Our method is evaluated on a range of hairstyles, including straight, wavy, curly, and coily hairs. To capture the appearance of short or coily hairs, our method comes with the support for hair cap and cross-card.
Hair; hair card; differentiable rendering
ccs: Computing methodologies Shape modeling
Our cards only          Our cards + textures          Strands model
Card #  200
200
200
800
100
351
100
Figure 1. Our automatic pipeline can convert a wide variety of strand-based hairstyles with 40K strands into hair card models using a small number of cards (bottom, depending on hairstyles) and 32 individual card textures, while preserving high visual fidelity. Here, we show the input strand model, our generated cards only, and cards rendered with textures.
[…]
1. Introduction
Hair plays a significant role in creating believable characters and immersive environments in films, video games, and virtual reality. However, an average human scalp has about 100,000 - 150,000 hair follicles. Although strand-based hair simulation and rendering have become increasingly popular in recent production pipelines due to their ability to realistically capture the dynamic behavior and visual richness (Tafuri, 2019; Epic Games, 2021; Huang et al., 2023; Hsu et al., 2023), hair cards remain a widely adopted technique in real-time applications (Jiang, 2016). Hair cards are flat, textured quad strips that approximate the visual appearance of hair clusters. As illustrated in Fig. 1, hair cards can effectively reproduce the intricate structure of complex hair styles. Beyond their role in representing high-fidelity hair with a large number of textured planes (Jiang, 2016), hair card representations are also widely used in level-of-detail (LoD), where strand-based models serve as a high-resolution representation, while hair cards can be used as a low-resolution approximation to significantly reduce rendering costs when hair is viewed from a distance.
Unfortunately, the industry currently heavily relies on crafting hairstyles manually using hair cards. In this workflow, artists first create a set of hair card textures that contain varying numbers of strands and degrees of curvature. They then manually extrude each hair card from the scalp outward, following the hair flow while varying its length, width, and orientation to achieve a realistic appearance. Apparently, the fidelity quality of the hair card model is primarily bound by the number of cards and textures allowed. As a result, manually crafting a low-resolution hair card model—with a limited number of quads and card textures while maintaining high visual fidelity to the reference strand-based model—is a challenging and labor-intensive task, typically requiring several days to weeks of work from an experienced artist.
Most recent research works have focused on reconstructing strand-based hair models from single-view images (Zhou et al., 2018; Zhang and Zheng, 2019; Wu et al., 2022; Zheng et al., 2023), multi-view inputs (Kuang et al., 2022; Wu et al., 2024b; Zhou et al., 2024), and CT scans (Shen et al., 2023). However, there is a lack of work on converting these reconstructed strands into hair card representations suitable for real-time applications. Extracting hair cards from strand-based hair models presents several challenges. First, there is a strict budget on both the number of hair cards and on the size of texture, imposed by memory and computational efficiency constraints. Therefore, the extraction process must carefully balance visual fidelity with performance. Second, the hairstyle is highly irregular and chaotic, showing significant variation in strand length, shape, and structure, making it challenging to generate a simplified representation while preserving visual details. Recently, Unreal Engine 5 (Epic Games, 2021) introduced a tool for the automatic generation of hair cards from strand-based models (Epic Games, 2025). However, the quality of its generated results remains insufficient for production use, often failing to preserve the appearance of the target hair strand model when the given hair card target is limited.
This paper presents an automated pipeline that converts strand-based hair models into hair cards. The process begins with hair strand clustering, where strands are grouped based on a hair shape similarity metric. For each cluster, we optimize the orientation of the card geometry to minimize the strand projection error. To improve memory efficiency and runtime performance, we perform second-stage clustering on the hair textures, enabling texture sharing across multiple cards. Finally, we jointly optimize the hair card positions, strand shapes, and hair widths to minimize differences in tangent, depth, and coverage between our output cards and the original strand-based model. To this end, conventional RGB-image texture representations can lead to significant aliasing artifacts. Instead, we propose an explicit hair card representation where each strand is projected into texture space as a 2D curve. These 2D curves serve as an intermediate representation that enables high-quality rendering and supports differentiable optimization. To enhance the visual fidelity for short and coily hairstyles, which is challenging for original hair card representation, our framework also supports baking a hair cap texture and generating a pair of crossed cards per hair cluster, enriching the volumetric appearance of the hair.
We evaluate our pipeline on a diverse range of hairstyles, including straight, wavy, curly, and coily hairs, with a limited number of cards and textures. Experiments show that our automated pipeline surpasses both UE automatic solution and manual-crafted cards.
2. RELATED WORK
This section briefly reviews work on hair representation, hair modeling, and extracting planar representations from meshes.
Hair Representation
Both hair simulation and rendering are computationally expensive due to the large number of individual strands and their complex interactions. To reduce computational costs, researchers have developed various reduced representations over the years, including 2D strips (commonly referred to as hair cards(Koh and Huang, 2001; Ward et al., 2003), cubic lattice structures(Volino and Magnenat-Thalmann, 2006), short hair strips (Guang and Zhiyong, 2002), and volumetric representations (Wu and Yuksel, 2016; Lee et al., 2019). During rendering, these reduced representations are expanded into full hair using baked textures or procedural functions. Hair cards remain widely used in the gaming industry due to their simplicity and efficiency (Jiang, 2016). For a comprehensive overview of hair rendering and simulation, we refer readers to the course by Bertails et al. (2008). However, creating hair cards for multiple LoDs continues to be a labor-intensive process for artists, representing a significant bottleneck in production pipelines.
(a) Strands
(b) Clustered strands
(c) Optimized cards
(d) Cards w/ exp. textures
(e) Texture reduction
(f) Final cards w/ textures
19.28/0.148/0.026
18.54/0.161/0.032
19.99/0.140/0.021
[…]
Hair Modeling
Traditionally, artists have used tools like Maya XGen to create strand-based and card-based hair models. However, manually crafting hair remains a labor-intensive process. To simplify hair authoring, Yuksel et al. (2009) introduced hair mesh, a volumetric representation that provides high-level editing tools for artists. Sketch-based input methods (Fu et al., 2007; Shen et al., 2020) have also offered user-friendly interaction capability, allowing artists to design hair directly on the screen. For automatic hair modeling, researchers have explored extracting strand-based geometries directly from images. Earlier methods relied on heuristic-based approaches (Kong and Nakajima, 1998; Paris et al., 2008; Jakob et al., 2009; Hu et al., 2017) or leveraged large 3D hair databases (Hu et al., 2015; Chai et al., 2016; Liang et al., 2018) for guidance. Recently, learning-based techniques have demonstrated accuracy and robustness in reconstructing simple hairstyles from single-view images (Zhou et al., 2018; Zhang and Zheng, 2019; Wu et al., 2022; Zheng et al., 2023) and multi-view inputs (Kuang et al., 2022; Rosu et al., 2022; Sklyarova et al., 2023; Wu et al., 2024b; Takimoto et al., 2024; Zakharov et al., 2024; Zhou et al., 2024). Wu et al. (2024a) present a geometric method to generate highly coiled hair. Beyond explicit modeling and reconstruction, hair synthesis methods enable hairstyle transfer from one to another using feature maps (Wang et al., 2009), further refined with learning-based features (Zhou et al., 2023; Sklyarova et al., 2024; Chen et al., 2024; He et al., 2025). However, these techniques primarily aim to generate explicit strand geometries, which are computationally intensive and unsuitable for real-time applications.
Billboard Extraction
In addition to hair, arbitrarily oriented billboards (or impostors) with textures are widely used for rendering trees and forests (Behrendt et al., 2005) and large-scale scenes (Lall et al., 2018; Hladky et al., 2022). To convert a 3D mesh to billboards, Décoret et al. (2003) proposed a greedy optimization algorithm that approximates the 3D model using a discretized plane parameterization in spherical coordinates. This method was later enhanced by Andujar et al. (2004) to produce better-fitting billboards. Silvennoinen et al. (2014) uses a similar concept to generate sets of planars as occluders. However, as hair presents unique challenges due to the extensive strand number and complex geometry, existing billboard generation methods cannot be directly applied.
3. Problem Statement and Overview
In this section, we first provide the problem statement, followed by an overview of our pipeline.
Problem Statement
The input to our method is a strand-based hairstyle model =$\mathcal{H}=\{\mathcal{S}_{i}\}$, where each strand $\mathcal{S}_{i}=\{\mathbf{s}_{i,j}\}$ is represented as a piecewise linear curve consisting of $n^{s}$ uniformly distributed samples $\mathbf{s}_{i,j}$ (with the same distance between two consecutive samples on each strand), along with a corresponding head mesh $\mathcal{M}^{\text{head}}$. The output of our pipeline is a hair card model $\mathcal{M}^{\text{output}}=\langle\{\mathcal{C}_{i}\}_{i=1}^{n^{c}},\{\mathcal{T}_{i}\}_{i=1}^{n^{t}}\rangle$, composed of $n^{c}$ hair cards and $n^{t}$ textures to balance visual fidelity and performance constraints. In order to save texture space, users oftentimes choose $n^{c}>\!\!>n^{t}$ such that multiple hair cards need to share the same texture. A hair card $\mathcal{C}_{i}$ refers to a quad strip with $n^{q}$ consecutive quads. Our objective is to ensure that the synthesized hair card model $\mathcal{M}^{\text{output}}$ approximates the visual appearance of the input $\mathcal{H}$ as closely as possible.
Overview
As illustrated in Fig. 2, our pipeline begins with hair strand clustering (Sec. 4.1), where strands are grouped into $n^{c}$ clusters based on a hair shape similarity metric. For each cluster, we then optimize the card orientation to minimize visual error, producing the initial card geometry (Sec. 4.2). Next, we introduce an explicit card texture representation by projecting the 3D curves onto a 2D texture space (Sec. 4.3). The generated hair textures are further clustered into $n^{t}$ groups to facilitate texture sharing across cards (Sec. 4.4). Finally, we jointly optimize the geometry and strands of all hair cards to minimize the difference between $\mathcal{H}$ and $\mathcal{M}^{\text{output}}$ using differentiable rendering (Sec. 4.5). The optimized cards and strands are then used to generate the final textures via rasterization.
4. Method
We revise the steps of our pipeline in this section.
4.1. Strand Clustering
Our method begins by clustering the input strands $\mathcal{H}$ into a number of subsets so each cluster can be represented as a hair card. We use a clustering method proposed by (Wang et al., 2009). Specifically, a hair shape similarity metric is defined as $\gamma(\mathcal{S}_{i},\mathcal{S}_{j})=\sum_{k=1}^{n^{s}}\|\mathbf{s}_{i,k}-\mathbf{s}_{j,k}\|_{2}^{2}/n^{s}$ to quantify the sample-wise distance between two hair strands $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$. Using this metric, we group $\mathcal{H}$ into clusters $\{\mathcal{G}_{k}\}$ with a k-mean clustering. Clearly, the center of each cluster is a mean strand denoted as $\bar{\mathcal{S}}_{k}=\{\bar{\mathbf{s}}_{k,j}\}$ with samples being $\bar{\mathbf{s}}_{k,j}=\sum_{\mathcal{S}_{i}\in\mathcal{G}_{k}}\mathbf{s}_{i,j}/|\mathcal{G}_{k}|$, where $\mathcal{G}_{k}\subseteq\mathcal{H}$ is a strand subset.
4.2. Card Geometry Initialization
Given the strand clusters ${\mathcal{G}_{k}}$, our next step is to initialize the hair card and prepare for further optimization. We first initialize an orthogonal frame along the mean strand and build initial card geometry. Then, we optimize orientation for each card to minimize projection error.
Card Geometry Fitting
To initialize the card geometry, we need to build an orthogonal frame system $\{\mathbf{t}_{k,j},\mathbf{n}_{k,j},\mathbf{b}_{k,j}\}$ at each sample $\bar{\mathbf{s}}_{k,j}$ along the central strand $\bar{\mathcal{S}}_{k}$, where $\mathbf{t}_{k,j},\mathbf{n}_{k,j},\mathbf{b}_{k,j}$ are the tangent, normal, and binormal vectors, respectively. The problem of building a frame system for a 3D curve has been thoroughly studied, for which the standard construction is the Frenet-Serret formulas. But this formula suffers from singular configurations and a more numerically stable choice is the Bishop formulas (Bergou et al., 2008), which have been adopted in strand-based hair simulation. Specifically, given the normal $\mathbf{n}_{k,1}$ for the first curve segment, the entire frame system can be computed by solving the Bishop formulas and we refer readers to (Bergou et al., 2008) for the piecewise linear discretization. Given the frame system, we calculate the maximum distance from $\bar{\mathbf{s}}_{k,j}$ to all corresponding strand samples within the cluster along the binormal direction such that:
$W_{k,j}=\max_{\mathcal{S}_{i}\in\mathcal{G}_{k}}\max_{\mathbf{s}_{i,j}\in\mathcal{S}_{i}}\left|(\mathbf{s}_{i,j}-\bar{\mathbf{s}}_{k,j})\cdot\mathbf{b}_{k,j}\right|.$
Figure 3. Example of hair cluster and its hair strip as well as per-sample frames.
Then, the card is constructed by connecting all $\mathbf{p}_{k,j}^{\pm}=\bar{\mathbf{s}}_{k,j}\pm W_{k,j}\mathbf{b}_{k,j}$ along the positive and negative $\mathbf{b}_{k,j}$ at each $\bar{\mathbf{s}}_{k,j}$ to form a quad strip with $n^{s}$ quads. Finally, we downsample the quad strip to use only $n^{q}$ quads, leading to our card geometry $\mathcal{C}_{k}$ as shown in Fig. 3.
Card Orientation Optimization
In the above discussion, we have assumed that the root normal $\mathbf{n}_{k,1}$ is given as the boundary condition of the Bishop formulas. This root normal $\mathbf{n}_{k,1}$ for the first card segment needs to be carefully computed to provide a good initial guess for further optimizations. We propose to optimize $\mathbf{n}_{k,1}$ by minimizing the difference between the strand cluster $\mathcal{G}_{k}$ and the hair card $\mathcal{C}_{k}$. To this end, we introduce the projection operator $\mathbf{s}_{i,j}^{\Gamma}=\text{argmin}_{\mathbf{s}\in\mathcal{C}_{k}}\|\mathbf{s}-\mathbf{s}_{i,j}\|$ as finding the closest point $\mathbf{s}$ on the quad strip of hair card $\mathcal{C}_{k}$, that is closest to $\mathbf{s}_{i,j}$. We then define the projection error as follows:
(1) $L^{\text{proj}}=\sum_{\mathcal{S}_{i}\in\mathcal{G}_{k}}\sum_{\mathbf{s}_{i,j}\in\mathcal{S}_{i}}\|\mathbf{s}_{i,j}^{\Gamma}-\mathbf{s}_{i,j}\|.$
Although the Bishop formulas are differentiable, the projection operator is non-differentiable, so gradient-based continuous optimization is non-available to minimize $L^{\text{proj}}$. Fortunately, the solution space is rather small. Since Eq. 1 is independent for different cards and due to the condition that $\mathbf{n}_{i,1}$ is orthogonal to $\mathbf{t}_{i,1}$, $\mathbf{n}_{i,1}$ essentially lies on a 3D circle. Therefore, we evenly sample a fixed number of potential $\mathbf{n}_{i,1}$ along the 3D circle and pick the direction leading to the smallest value of Eq. 1. At this point, we have initialized all the card geometries.
4.3. Explicit Hair Card Texture
In the standard rendering pipeline, textures are introduced for each hair card to represent various attributes in the form of RGB images. However, the hair strands are extremely thin, and rasterizing them into RGB textures would introduce severe aliasing error and noisy gradient information. To mitigate this issue, we propose a novel intermediary explicit hair card representation. Specifically, each hair strand $\mathcal{S}_{i}\in\mathcal{G}_{k}$ is first projected onto the quad strip $\mathcal{C}_{k}$. Specifically, for each $\mathbf{s}_{i,j}\in\mathcal{S}_{i}$, we find the closest point $\mathbf{s}_{i,j}^{\Gamma}$ lying on the associated quad strips $\mathcal{C}_{k}$. The corresponding 2D uv-coordinate $\mathbf{s}_{i,j}^{\text{uv}}=(u_{i,j},v_{i,j})\in[0,1]^{2}$ on the texture can be computed such that the 3D world position of each sample $\mathbf{s}_{i,j}$ can be reconstructed as:
(2) $\mathbf{s}_{i,j}=u_{i,j}\mathbf{p}_{i,j}^{0}+v_{i,j}\mathbf{p}_{i,j}^{1}+(1-u_{i,j}-v_{i,j})\mathbf{p}_{i,j}^{2}+z\mathbf{n}_{i,j}^{\star},$
where $\mathbf{p}_{i,j}^{\bullet}$ are the vertices of the corresponding triangle face in the card geometry that contains $\mathbf{s}_{i,j}^{\Gamma}$, $\mathbf{n}_{i,j}^{\star}$ is the normal of that triangle, and $z\in\mathbb{R}$ is the displacement along $\mathbf{n}_{i,j}^{\star}$. Hence, each hair card texture can be explicitly represented as a set of points $\{\mathbf{s}_{i,j}^{\text{uv}}\}$ embedded within the uv-space. Using $\{\mathbf{s}_{i,j}^{\text{uv}}\}$ as our intermediary representation induces two remarkable features. First, this representation is amenable to end-to-end differentiable rendering, while it does not suffer from aliasing error induced by RGB images. Indeed, for differentiable rendering, 3D hair strands can be first recovered from Eq. 2. These 3D strands are then extended as camera-facing triangle strips with a pre-strand hair width, denoted as $w_{i}$, which is then forwarded to differentiable renderer. Each and every step of this procedure is differentiable and a similar technique is proposed in (Sklyarova et al., 2023). Further, our representation can be converted back to RGB images via rasterization. Finally, to avoid uv-space tangent distortion by the card geometry in world space, we decompose tangent from strand geometry and store it as a per-vertex 3D attribute denoted as $\{\mathbf{t}_{i,j}^{\text{3D}}\}$, which plays a central role in the hair appearance model.
4.4. Texture Reduction
To facilitate texture sharing and improve both memory efficiency and runtime performance, user could optionally reduce the number of textures before packing them into a texture atlas. Clusters with similar strand counts and shapes are grouped to share the same texture. Specifically, after the per-cluster optimization, we allow users to provide the target texture number $n^{t}$. We then perform another round of k-mean clustering to merge cards with similar textures. To this end, we first rasterize the uv-space strands into RGB textures and then compute the Learned Perceptual Image Patch Similarity (LPIPS) metric (Zhang et al., 2018) between all pairs of card textures. The LPIPS metric is used to guide our k-mean clustering. For each cluster, only the texture closest to the k-mean center is retained, and it is reused across all hair cards within the same cluster. After the clustering, our joint optimization would then fine-tune the shared textures to collectively match the appearance.
4.5. Joint Optimization
After texture reduction, we perform another optimization with the visual losses and geometric regularization to match the collective visual appearance between the input strands =$\mathcal{H}=\{\mathcal{S}_{i}\}$ and our final cards $\{\mathcal{C}_{i}\}$ with uv-space strands $\{\mathbf{s}_{i,j}^{\text{uv}}\}$ and tangent $\{\mathbf{t}_{i,j}^{\text{3D}}\}$.
Optimization
Given the differentiable renderer, we fine-tune both the card geometry and strand textures for each cluster using gradient-based optimization by solving the following optimization:
(3) $\underset{\mathbf{p}_{k,j}^{\pm},~{}\mathbf{s}_{i,j}^{\text{uv}},~{}\mathbf{t}_{i,j}^{\text{3D}},~{}w_{i}}{\text{argmin}}\quad\sum_{v}\lambda^{v}L^{v}+\sum_{g}\lambda^{g}L^{g},$
[…]
Visual Losses
While our ultimate goal is to produce a hair card model that looks identical to the input strand model, rendering complicated light interactions between a large number of hair segments is computationally intractable. In particular, the interaction between light and a hair strand is described by the Bidirectional Curve Scattering Distribution Function (BCSDF) denoted $f(\theta_{i},\phi_{i},\theta_{o},\phi_{o})$, as proposed by Marschner et al. (2003). Here, $\langle\theta_{i},\phi_{i}\rangle$ and $\langle\theta_{o},\phi_{o}\rangle$ represent the incoming and outgoing light directions, respectively. The inclination angle $\theta_{\bullet}$ represents the deviation from the strand normal plane, while the azimuth angle $\phi_{\bullet}$ captures the orientation around the hair axis, both derived from the strand tangent. In order for more efficient optimization, instead of matching the final appearance, our optimization matches the tangent channel for shading and the depth channel for strand position along the view direction. In particular, tangent and depth losses are measured by the Mean-Squared Error ($\mathsf{MSE}$) over all views as:
[…]
where $\mathcal{I}^{v}(\bullet,\omega)$ is the rendering function that rasterizes the channel ${v\in\{\text{tangent},~{}\text{depth}\}}$ of entities $\bullet$ under view direction $\omega$. We further add the following dice loss (Sudre et al., 2017) for matching the silhouette:
(5) $L^{\text{dice}}=\int_{\mathbb{S}^{2}}1-D\bigl{(}\mathcal{I}^{\text{mask}}(\mathcal{H},\omega),\mathcal{I}^{\text{mask}}(\langle\mathcal{C}_{k},\mathbf{s}_{i,j}^{\text{uv}},w_{i}\rangle,\omega)\bigr{)}d\omega,$
where $D(A,B)={2|A\cap B|}/{(|A|+|B|)}$ is the Sørensen-Dice coefficient to quantify the similarity between two masks, $A$ and $B$, where $|A|$ and $|B|$ are the total number of pixels in $A$ and $B$, and $|A\cap B|$ is the common area between $A$ and $B$. For rendering the binary mask image, we set all pixels to be one in a all-zero background. Note that we do not need the tangents $\mathbf{t}_{i,j}^{\text{3D}}$ for mask rendering.
Geometric Regularization
For accurate visual appearance, we treat $\mathbf{t}^{\text{3D}}_{i,j}$ as a separate decision variable that is not bound to the strand geometry. However, we still encourage the geometric definition of tangent is consistent with the optimized tangent direction, via the regularization:
(6) $L^{\text{match}}=\sum_{\mathcal{G}_{k}}\sum_{\mathcal{S}_{i}\in\mathcal{G}_{k}}\sum_{j=1}^{N^{s}}\left\|\mathbf{t}^{\text{3D}}_{i,j}-\frac{\mathbf{s}_{i,j+1}-\mathbf{s}_{i,j}}{\|\mathbf{s}_{i,j+1}-\mathbf{s}_{i,j}\|}\right\|^{2}.$
Finally, we introduce a collision loss to prevent hair cards from penetrating the head mesh, formulated as:
(7) $L^{\text{collision}}=\sum_{\mathcal{G}_{k}}\sum_{j=1}^{N^{s}}\sum_{\bullet\in\{+,-\}}\|\min(0,\text{SDF}(\mathbf{p}_{k,j}^{\bullet},\mathcal{M}^{\text{head}}))\|^{2},$
where SDF is the signed distance from a point to the head mesh.
Texture Baking
As our final step, after optimization, we rasterize the UV-space strands and tangents into RGB-format textures to generate the final tangent, depth, and alpha maps, as shown in Fig. 4. Additionally, we bake an ambient occlusion (AO) texture to enhance visual detail and store local lighting information on the hair cards. Note that the $z$ offset along normal direction in Eq. 2 is not optimized, which is essential for accurate AO baking.
Tangent
Depth
Alpha
AO
Figure 4. Example of our output hair card textures
[…]
4.6. Extensions
To enhance both the visual fidelity of our output and the practical usability of our pipeline, we introduce the following two additional features, which can be tuned on optionally.
Crossed Cards
To simulate hair volume and improve visual fidelity from multiple viewing angles, we introduce crossed hair cards during generation. Specifically, for each cluster, we generate a pair of hair cards placed at a 90-degree angle to each other, creating the illusion of volumetric hair using flat geometry, as shown in Fig. 5. In practice, during the card generation stage (Sec. 4.2), we simply create an additional card perpendicular to the primary card at each segment by rotating both $\mathbf{n}_{k,j}$ and $\mathbf{t}_{k,j}$ over 90 degrees, giving a crossed configuration.
Strands
UE cards
Our cards
Our crossed
Figure 5. Compared with single cards from Unreal Engine auto tool (Epic Games, 2025) and ours, our crossed cards can create a volumetric appearance, especially around the ponytail.
[…]
Strands
Our crossed
Our cap
Our crossed + cap
Figure 6. Our hair cap can improve the scalp coverage and hairline.
[…]
Hair Cap
Since hair cards alone may not fully cover the scalp under certain camera angles or lighting conditions, particularly near the roots, we construct a hair cap, a base layer of geometry and texture applied to the scalp that simulates the appearance of short, dense hair, so, it can efficiently represent root fuzz and base hair density with minimal computational cost. To create the hair cap mesh and its associated texture, we begin by projecting all hair roots onto their nearest faces on the head mesh $\mathcal{M}^{\text{head}}$. We then extract all faces containing at least one hair root, along with their one-ring neighboring faces to form the hair cap mesh. In order to prevent z-fighting during rendering, we slightly extrude the vertices of the cap mesh along their normal directions by a small distance $\epsilon^{\text{cap}}$. Finally, for each strand, we extract the first sub-segment with a fixed arc-length of $\epsilon^{\text{root}}$ and bake the sub-segment onto the scalp texture. Specifically, we bake the tangent, alpha, and AO into three texture channels as shown in Fig. 6.
5. Results
We conduct all experiments on a system equipped with an AMD Ryzen Threadripper 3970X 32-core CPU, 256 GB of memory, and an NVIDIA RTX 3090 GPU. Our framework is implemented in PyTorch and customizes Nvdiffrast (Laine et al., 2020) as our backbone differentiable render.
We evaluate our pipeline on a diverse set of hairstyles, including short, bangs, straight, ponytail, bun, wavy, curly, and coily, from MetaHuman (Epic Games, 2021) and dataset in previous high-fidelity hair reconstruction work (Shen et al., 2023). All hair models are down-sampled to use 40K hair strands, each with $n^{s}=32$ samples. Following the hair card texture settings used in MetaHuman, we set our output texture atlas size to $2048\times 2048$, which accommodates $n^{t}=4\times 8=32$ individual hair card textures, each at a resolution of $512\times 256$ by default. The card orientation optimization uses a screen size $128\times 128$, while joint optimization uses $256\times 256$ to capture strand-level details. To assess the visual similarity to the input strand-based model, we use three averaged metrics, PSNR $\uparrow$ for the shading evaluation, LPIPS $\downarrow$ for perceptual similarity, especially high-frequency hair details, and coverage error $\downarrow$ for the hair silhouette. The average values of these metrics are computed over 12 uniformly distributed viewpoints around the hair.
Ablation study on # card and # texture
Fig. 11 presents a comparison between our method and Unreal Engine’s automatic hair card generator (Epic Games, 2025) under varying number of card and texture budgets. We evaluate results using 200, 100, and 50 cards, targeting low-resolution LoD hair generation. For textures, we test with 16 and 32 textures, following the industrial best practices for texture budgeting. As expected, visual similarity to the reference strand-based model decreases as the card and texture budgets are reduced. Therefore, key to creating a good hair model lies in finding a balance between visual quality and geometric efficiency. Nevertheless, our method consistently achieves higher visual fidelity in all three metrics, even when using fewer textures and only 50 cards.
Ablation study on losses
We further conduct an ablation study on all loss and regularization terms on a straight hairstyle from MetaHuman (Epic Games, 2021) with 100 hair cards and 32 textures. As shown in Fig. 7, each component is necessary to achieve consistent and accurate results. For the rest of experiments, we adopt two sets of loss weights for different hair types. For straight hair, we use $\lambda^{\text{tangent}}=10$, $\lambda^{\text{depth}}=10$, $\lambda^{\text{dice}}=5$, $\lambda^{\text{match}}=3$, and $\lambda^{\text{collision}}=1\times 10^{5}$. For curly hair, we use a different configuration as $\lambda^{\text{tangent}}=5$, $\lambda^{\text{depth}}=15$, $\lambda^{\text{dice}}=3$, $\lambda^{\text{match}}=3$, and $\lambda^{\text{collision}}=1\times 10^{5}$, to encourage more curl strands.
19.83
19.69
19.71
Strands
Our total
w/o tangent loss
w/o depth loss
19.82
19.67
19.53
w/o collision loss
w/o dice loss
w/o match loss
[…]
Comparison with optimizing texture directly
We validate the necessity of our explicit strand representation by comparing it to a conventional differentiable rendering pipeline that directly optimizes hair card textures using Nvdiffrast (Laine et al., 2020). In this experiment, we provide a fixed hair card geometry along with tangent images from eight view directions. We then use Nvdiffrast to optimize the texture mapped to the card so that its rendered appearance matches the given tangent images as closely as possible from those views. Although Nvdiffrast can produce hair textures with distinct shape features when optimizing for a single view direction, extending the optimization to multiple views leads to noisy and inconsistent results. Introducing a smoothness loss can eliminate noise at the cost of losing high-frequency details, as shown in Fig. 8. More importantly, this texture-based approach lacks the spatial structure for computing ambient occlusion accurately, as hair-to-hair spatial relationships are not preserved during texture optimization.
Strands
Input card
Tex-based
w/o smooth
Tex-based
w/ smooth
Ours
Figure 8. Given input card and hair views from eight directions, directly optimizing texture leads to noisy results, while adding smoothness loss eliminates high-frequency hair structure. Our strand-based optimization avoids the above issues effectively.
[…]
Comparison with Unreal Engine’s auto-generated and manual-crafted cards
We compare our results with the output of Unreal Engine’s built-in automatic hair card generation tool (Epic Games, 2025) and artist-crafted hair cards from MetaHuman (Epic Games, 2021). We compare with the lowest LoD hair card model from MetaHuman, which only contains 351 cards and 20 textures. As shown in Fig. 9, using the same set of 351 hair cards and 20 textures, our method achieves better visual fidelity to the input strand-based model than both Unreal Engine’s auto-generated and manual-crafted cards. While our higher PSNR partly benefits from more accurate AO textures, as we preserve the displacement during optimization, the lower LPIPS and coverage errors clearly demonstrate that our results more faithfully preserve hair occupancy and shape.
Strands
Manual cards
UE auto cards
Our cards
17.71/0.279/0.044
17.78/0.209/0.038
19.99/0.140/0.021
Figure 9. Comparison with Unreal Engine’s auto-generated and manual-crafted cards. Our method outperforms the other two in all metrics with the same 351 cards and 32 textures.
[…]
Cards extraction performance
Given input hair with 40K strands with $n^{s}=32$ samples, the average process time of our card generation pipeline for cards with 100 cards and 32 textures is about 46 minutes, where joint optimization takes about $42\%$ of the computation time. For a larger configuration of 400 cards with the same number of textures, the total runtime increases to approximately 1 hour, with joint optimization comprising roughly one-third of the computation time. In both cases, the optimization process converges within roughly $200$ epochs. Memory cost is about 7 GB for $100$ cards and 10 GB for $400$ cards, respectively.
Table 1. Statistics about card # used in both our method and Unreal Engine’s auto cards, as well as the corresponding visual metric statistics, PSNR $\uparrow$, LPIPS $\downarrow$, and coverage error $\downarrow$. Crossed card is indicated by ×$\times 2$.
Unreal Engine (Epic Games, 2025) Ours
Model Fig. # Card # PSNR $\uparrow$ LPIPS $\downarrow$ Coverage $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ Coverage $\downarrow$
Ponytail 5 100$\times$2 20.91 0.123 0.022 22.70 0.095 0.010
Short 6 100$\times$2 22.34 0.112 0.014 23.51 0.100 0.010
Bun 9 351 17.78 0.209 0.038 19.99 0.140 0.021
Straight 11 100 17.39 0.230 0.038 18.21 0.176 0.034
Curly 12 400 18.54 0.153 0.038 20.89 0.138 0.016
Bangs 12 100 18.09 0.218 0.032 20.47 0.154 0.018
Blowout 12 200 17.16 0.200 0.038 20.44 0.162 0.017
Wavy 12 200 16.82 0.228 0.041 19.89 0.166 0.016
Coily 13 400$\times$2 17.90 0.174 0.032 19.09 0.131 0.022
Real-time demo
Fig. 10 demonstrates that our hair cards can be used directly in Unreal Engine with real-time rendering and simulation. Specifically, we generate one guide hair per card for simulation and drive the deformation of the hair cards using linear blend skinning. Simulating a hair card model with 512 cards takes only 0.12 ms per frame. Please refer to the supplemental video.
Figure 10. Simulating our output hair card model in Unreal Engine, taking only 0.12 ms per frame.
[…]
6. Conclusion
We have presented a fully automated pipeline for converting strand-based hair models into efficient and visually compelling hair card representations. By leveraging a differentiable rendering framework, our method first clusters the hair strands to initialize card geometry, and then clusters the hair textures for sharing to reduce the memory cost. Finally, we conduct jointly optimization over both card geometry and textures. A key contribution of our approach is the introduction of 2D curve-based texture encoding, which offers a resolution-independent representation that effectively captures fine strand details while remaining compatible with differentiable rendering. Our method supports a wide range of hairstyles and lengths, and introduces mechanisms such as hair caps and crossed card generation to handle visually complex hair types like short and coily styles. Additionally, the ability to share textures across LoDs makes our approach well-suited for real-time applications where memory efficiency and performance are critical.
Limitations
While our pipeline demonstrates strong performance and visual fidelity across a range of hairstyles, several limitations remain. First, the optimization is conducted in multiple substeps rather than as a fully end-to-end differentiable pipeline, which may limit global consistency. Second, the loss functions primarily rely on visible attributes such as tangent alignment, depth, and mask coverage. However, for complex hairstyles, especially those with dense or layered structures, interior strand arrangements play an important visual role and are not explicitly enforced. Third, our method assumes a uniform hair color across the entire hairstyle, which may not generalize well to stylized or multi-colored hair. Finally, although our framework supports a wide range of hair types, coily hair remains a challenge due to its highly curved geometry, as shown in Fig. 13, which typically requires a large number of cards for faithful representation.